OpenStudy (anonymous):

The list price for a lawnmower is $420. The net price is $315. What is the trade discount for the lawnmower? I am getting 25%, can someone explain to me why I keep getting it wrong because I'm doing it the exact same way I've done the other problems and this one keeps telling me I'm wrong

OpenStudy (anonymous):

discount=420-315=? discount %\[=\frac{ discount }{ list~ price }\times 100\]

OpenStudy (anonymous):

That comes out to 25%.... This question has me stumped though I'm not undertsanding where the trade discount part is coming from

OpenStudy (anonymous):

How is it a trade?

OpenStudy (anonymous):

Maybe its 105 dollars

OpenStudy (anonymous):

Since its not asking for the percentage but the actual discount

OpenStudy (anonymous):

Yeah you're right, I guess I misunderstood the actual question, thank you

OpenStudy (anonymous):

Np